find the equation that represents a line l that goes through (–2, –9) and has a slope of 5
Answer by oscargut(2103)

line l that goes through (–2, –9) and has a slope of 5
in gral if P=(x0,y0) and slope=m
line is:
y-y0=m(x-x0)
so equation of the line l is:
y+9=5(x+2)
or
y=5x+1